Louise is twice as old as her younger brother John. Their combined ages are 45. How old is each person?
Found 2 solutions by Nieleya, unlockmath:
Answer by Nieleya(22)

Louise is twice as old as her younger brother John. Their combined ages are 45. How old is each person?
2a + 1a = 45
3a = 45
a = 45/3
a = 15
Louise is 30 years old and John is 15 years old..

Answer by unlockmath(1688)

Hello, We can set "x" to represent John's age and 2x for Louise age. Now we can set up an equation to look like this:
x + 2x = 45 Combine like terms to be:
3x=45 Divide both sides by 3 to give us:
x=15
Now we know how old John is and from here I'll let you figure out Louise's age.
RJ
Check out a new book I wrote at:
www.math-unlock.com